OpenStudy (anonymous):

Solve for u. -17 + 2u = -u + 19

pooja195 (pooja195):

Welcome to openstudy :) Any ideas on what the first step maybe?

OpenStudy (arihangdu):

2u+u=19+17 3u=46 U=46/3 U=15 1/3

OpenStudy (arihangdu):

Hope this will help you?